From GlueXWiki
Jump to: navigation, search

August 23, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Cathodes
    • Third package
  2. Cathode corrosion
    • Results from SEM analyses: [1]
    • Test chambers
  3. Engineering (Dave)
    • Cathode strong-backs
  4. Electronics (Chris, Nick)
    • PCBs
  5. TDR, pages 132-145 (Lubomir)
  6. Other


Participants: Fernando, Tim, Eugene, Dave, Simon, Vlad, and Lubomir


- Dave: working on 10 new cathodes to be used in the first package, not in the second package as we always stated before. Again the bottleneck is the card gluing, but we will start making progress there, by tomorrow we plan to have the cathode strong-backs ready to be used to move the cathodes. Will start working soon on the other two cathodes that are needed to finish the first package. After that we will have two sets of cathode foils left.

- The third package is ready with the subsystems: grounding, cooling tubes, HV cabling. The plan is to move it to 126 tomorrow. The grounding worked fine, Vlad put ~100 clips for grounding as planned and another ~100 attached close to both sides of the cards (where possible and needed) to prevent the cathode deformation at the periphery. As Casey suggested, we used a new method to install the cooling tubes. We tightened the bracket on one card in the middle of the tube and installed the rest of the brackets loose. Then, going from the middle card toward the two ends of the tube we tightened the rest of the brackets. Thus, the copper tubes apply some pressure on the cards; before we had problems with some of the cards being pulled by the tubes. Gluing of the back brackets with green epoxy to the tubes was done at the end. In 126 after installing the cables we should experiment how easy will be to remove the extra clips we put on the card sides.

Cathode corrosion

- New SEM results from Olga (follow the above link) for the samples from the Vlad's test with the vacuum chamber; the chamber had water inside and the temperature at end of the day (lamps were off overnight) was close to 100degC. The test was done for about 4 days:

  • Copper foil with EPDM inside: significant corrosion, a lot of sulfur present.
  • With the EPDM of the gasket, part of the foil strips were inside the gas chamber and part exposed to the air: no sulfur seen at the air side and a lot of it at the gas side. However, inside the gas volume there was moisture and the temperature was much higher than outside. We had seen corrosion before also in the samples placed in air with some water put on the foils and no heating. We will do more tests.
  • Copper foil with Viton: some traces visible, almost no sulfur there but some chlorine. Tim suggested cleaning the O-rings before installation.
  • With Viton and Apiezon: no visible corrosion, only copper and carbon lines are visible, but the grease was not cleaned. Olga will repeat the measurements after cleaning it.

- The first testing chamber with production wire frame and cathodes was opened yesterday, exactly 4 weeks operating with HV and gas. Some traces from the EPDM O-ring are visible but it doesn't look like the corrosion on the first and second packages. We conclude that the humidity is an important factor. No traces on the cathode against the Viton+Apiezon O-ring are visible. The test with the small-scale prototype is running two weeks, haven't opened it yet.

- Tim for the humidity conditions in the Hall: nominally it will be less than 45%, but he can't exclude some periods when it might be higher. There will be dry air blowing through the magnet. We should avoid opening the Hall doors for a long time during the installation.


- We will have two strong-backs to be used to grab the cathodes from both sides and keep them flat. We discussed how to support the cathode periphery when using the strong-backs. Casey made a table with rods that will be used to store the cathodes in a flat position.

- The rings for the cathode strong-backs came from the REX company right after the meeting, but there were issues with the fittings for the suction cups; also the weight of the aluminum ring seams too big.


- Fernando: wrong zip file was send to the company for the PCBs; now all the files have the correct names to exclude such a problem.

- Main discussion was what to do the with these four PCB sets. Mechanically, compared to the original PCBs, there is a small difference in the gas connector positions (about a mm); according to Tim it will work and mechanically the wire planes will be interchangeable. Fernando: some of the signal traces are at different positions and the gold-plated ground area is bigger; the traces going through that area have to be covered with conformal coating. There should be no differences in the HV PCBs, but Fernando will check all the boards and on Monday we will decide if these can be used. Eugene: if you spend more than two weeks on these boards we better purchase new ones, which however will be tricky (at the end of the fiscal year) and will delay the production. Lubomir and Dave: we will need the boards in two weeks, if not we have to re-work the schedule.


- Lubomir received corrections from Bill and Simon, still some work remains to be done here.


- Eugene urges us to do all the purchases for the FDC as soon as possible at least by the beginning of September; after that most likely there will be no money till the end of FY12.